# Complexity results and active-set identification of a derivative-free method for bound-constrained problems

### v2

@@ 1 Introduction
In the current paper, we analyze a line search method to solve problem (1). The algorithm under analysis is a modification of the one proposed in [29], equipped with a line search technique described in [7]. In particular, for any considered direction (i.e., a vector of the canonical basis), the line search technique first checks for a sufficient decrease in the objective function using a giving stepsize. Then, if such a decrease is obtained, an extrapolation (or expansion) phase starts, where increasingly larger values of the stespize are tried until some conditions are met. This approach allows us to obtain complexity and identification results that extend those existing in the literature for direct-search [19] and model-based [22] methods. In particular, let us summarize the main contributions of this paper below.
@@
    Here, we show that the proposed algorithm correctly identifies the active constraints satisfying the strict complementarity condition in a finite number of iterations, without using any parameter-dependent estimate. Namely, this feature is just an intrinsic property of the proposed algorithm. Such a result is obtained by using some tools from the analysis of derivative-based methods  [13, 32]. More specifically, we define a measure which represents the minimum strict complementarity among the active constraints, thereby providing a neighborhood of the limit points where the active-set identification holds. Let us remark that also this identification result is obtained thanks to the extrapolation techninque used in the line search procedure, which allows the stepsize to expand until we hit the border of the feasible set when we are in a neighborood of a stationary point.
@@ 2 The algorithm
This section is concerned with the definition of a line search algorithm to solve problem (1). The proposed method, denoted as Algorithm 1, is a slightly modified version of the one proposed in  [28] and uses some ideas from the one proposed in [7] for unconstrained problems.
@@
Given a feasible point $x$, the exploration of the $i$th coordinate direction $e_{i}$ is performed by a line search procedure outlined in Algorithm 2. First, we check if the given stepsize $\nu$ is feasible along ±$\pm e_{i}$, that is, if one between $x+\nu e_{i}$ and $x-\nu e_{i}$ is feasible. If this is not the case, then we quit the line search returning a zero step length to indicate a failure. Otherwise, we try to determine if one between $e_{i}$ and $-e_{i}$ is a “good” descent direction, that is, if a sufficient decrease in the objective function can be obtained by using a feasible stepsize. If neither $e_{i}$ nor $-e_{i}$ qualifies as a suitable descent direction, then the line search procedure terminates, returning a zero step length to indicate a failure. Conversely, if a sufficient decrease of $f$ is obtained, then an extrapolation (or expansion) phase starts (i.e., lines 12–15), where we try to increase the stepsize to the maximum extent while preserving feasibility and guaranteeing the sufficient decrease condition. Specifically, the while loop keeps expanding the stepsize as long as the most recently accepted point remains strictly within the bounds (i.e., $\alpha<\alpha_{\text{max}}$) and the new tentative point is sufficiently better than the last accepted one (i.e., $f(x+\eta d)\leq f(x+\alpha d)-\gamma(\eta-\alpha)^{2}$).
@@
12:set $\alpha=\bar{\alpha}$ and $\eta=\min\{\alpha/\delta,\alpha_{\text{max}}\}$
13:while ($\alpha<\alpha_{\text{max}}$ and $f(x+\eta d)\leq f(x+\alpha d)-\gamma(\eta-\alpha)^{2}$) do
14:     set $\alpha=\eta$ and $\eta=\min\{\alpha/\delta,\alpha_{\text{max}}\}$
@@ 3.1 Assumptions and preliminary results
The above measure has been successfully used in the analysis of some direct search methods for linearly constrained problems [19, 23, 24]. It can be interpreted as the progress on a first-order model in a ball centered at $x$ with unit ray subject to feasibility constraints [11, 24], thus generalizing $\|\nabla f(x)\|$ from the unconstrained setting. Originally proposed in [10] for more general constraints and further analyzed in [11], $\chi(x)$ is continuous, non-negative and such that $\chi(x)=0$ if and only if $x$ is a KKT point. So, we can define a stationary point as follows.
@@
$\Delta_{k+1}=\tilde{\alpha}_{k+1}^{\bar{\imath}}=\theta\Delta_{k}.$
@@ 3.2 Global convergence
$f(x+\eta d)\leq f(x+\alpha d)-\gamma(\eta-\alpha)^{2},$ (6)
with $\eta=\alpha/\delta=\bar{\alpha}/\delta^{j+1}$. Now, for $j$ sufficiently large, (6) contradicts Assumption 1, i.e., that $f$ is bounded from below on the feasible set. ∎
In the following, we present some results concerning the global convergence of Algorithm 1 to stationary points. More specifically, by extending some results from [7, 28], we establish a relationship between $\nabla f(x_{k})$, with specific directions $d$, and the largest tentative step length at $\Delta_{k+1}$. As to be shown, the bound depends, besides on the problem dimension $n$, on the Lipschitz constant $L$ and the algorithm parameters $\gamma$, $\theta$ and $\delta$.
@@
First, consider an iteration $k$ such that $x_{k+1}\neq x_{k}$ (i.e., a successful iteration). The following cases can occur, recalling that the analysis is limited to considering directions ±$\pm e_{i}$, $i=1,\ldots,n$, belonging to $G_{k}$.
@@
    $e_{i}\in G_{k}$, $y_{k}^{i+1}=l_{i}$ and $y_{k}^{i}=l_{i}$. Then, $\alpha_{k}^{i}=0$ and $\tilde{\alpha}_{k+1}^{i}=\nu_{k}^{i}$. From the instructions of the line search procedure, we have that
@@
    $y_{k}^{i+1}=l_{i}$ and $y_{k}^{i}>l_{i}$. Then, $\tilde{\alpha}_{k+1}^{i}=\alpha_{k}^{i}>0$ and $-e_{i}\in G_{k}$. From the instructions of the line search procedure, there exists $\beta_{k}^{i}\in\{0,\delta\tilde{\alpha}_{k+1}^{i}\}$ such that
@@
    Hence, by similar reasonings as in case (1a), we obtain
@@
    $-e_{i}\in G_{k}$, $y_{k}^{i+1}=u_{i}$ and $y_{k}^{i}=u_{i}$. Then, $\alpha_{k}^{i}=0$ and $\tilde{\alpha}_{k+1}^{i}=\nu_{k}^{i}$. Reasoning as in case (1a), with minor differences, we get
@@
    $y_{k}^{i+1}=u_{i}$ and $y_{k}^{i}<u_{i}$. Then, $\tilde{\alpha}_{k+1}^{i}=\alpha_{k}^{i}>0$ and $e_{i}\in G_{k}$. Reasoning as in case (1b), with minor differences, we get
@@
    {±$\{\pm e_{i}\}\cap G_{k}\neq\emptyset$, $l_{i}<y_{k}^{i+1}<u_{i}$ and $y_{k}^{i}=y_{k}^{i+1}$. Then, $\alpha_{k}^{i}=0$ and $\tilde{\alpha}_{k+1}^{i}=\nu_{k}^{i}$. From the instructions of the line search procedure, we have that
@@
    $l_{i}<y_{k}^{i+1}<u_{i}$ and $y_{k}^{i}\neq y_{k}^{i+1}$. Then, $\alpha_{k}^{i}>0$, $\tilde{\alpha}_{k+1}^{i}=\alpha_{k}^{i}$ and $d_{k}^{i}\in G_{k}$. Let us assume that $d_{k}^{i}=-e_{i}$, i.e., $y_{k}^{i+1}=y_{k}^{i}-\tilde{\alpha}_{k+1}^{i}e_{i}$ (the proof for the case $d_{k}^{i}=e_{i}$ is identical, except for minor changes). Then, from the instructions of the line search procedure, there exist $\beta_{k}^{i}$ and $\bar{\beta}_{k}^{i}$ such that $\beta_{k}^{i}\in\{0,\delta\tilde{\alpha}_{k+1}^{i}\}$, $\tilde{\alpha}_{k+1}^{i}<\bar{\beta}_{k}^{i}\leq\tilde{\alpha}_{k+1}^{i}/\delta$ and
@@
$\chi(x_{k})\leq\sqrt{n}\left(-\nabla f(x_{k})^{\top}d+\frac{\Delta_{k+1}}{\theta}\right)M_{g}.$
@@
$\Phi_{k}=f(x_{k})+\frac{1}{2}c^{2}\gamma\Delta_{k}^{2}.$ (16)
@@
$c_{1}=\gamma\min\left\{\left(1-\frac{c^{2}}{2}\right),\frac{c^{2}}{2},\frac{1}{2}c^{2}\left(\frac{1-\theta^{2}}{\theta^{2}}\right)\right\}>0,$ (19)
@@
  • $x_{k}\neq x_{k-1}$ (i.e., $k-1$ is a successful iteration) and $\Delta_{k}=\Delta_{k-1}$. Hence, there exists a coordinate ${\bar{\imath}}$ such that we have moved from $x_{k-1}$ along ±$\pm e_{\bar{\imath}}$ and we have
@@
    $\begin{split}\Phi_{k}-\Phi_{k-1}&=f(x_{k})-f(x_{k-1})+\frac{1}{2}c^{2}\gamma\left(\Delta_{k}^{2}-\Delta_{k-1}^{2}\right)\\ &\leq-\gamma c^{2}\Delta_{k}^{2}+\frac{1}{2}c^{2}\gamma\left(\Delta_{k}^{2}-\Delta_{k-1}^{2}\right)\\ &\leq-\gamma c^{2}\Delta_{k}^{2}+\frac{1}{2}c^{2}\gamma\Delta_{k}^{2}\\ &=-\gamma\frac{c^{2}}{2}\Delta_{k}^{2}.\end{split}$ (20)
@@
    Then, from the line search procedure, we can write
    $f(x_{k})\leq f(x_{k-1})-\gamma\Delta_{k}^{2}.$
@@
    $\begin{split}\Phi_{k}-\Phi_{k-1}&=f(x_{k})-f(x_{k-1})+\frac{1}{2}c^{2}\gamma\left(\Delta_{k}^{2}-\Delta_{k-1}^{2}\right)\\ &\leq-\gamma\Delta_{k}^{2}+\frac{1}{2}c^{2}\gamma\left(\Delta_{k}^{2}-\Delta_{k-1}^{2}\right)\\ &\leq-\gamma\Delta_{k}^{2}+\frac{1}{2}c^{2}\gamma\Delta_{k}^{2}\\ &=-\gamma\left(1-\frac{c^{2}}{2}\right)\Delta_{k}^{2}.\end{split}$ (21)
  • $x_{k}=x_{k-1}$ (i.e., $k-1$ is an unsuccessful iteration). Hence, for all $i\in\{1,\dots,n\}$ we have either
    $\tilde{\alpha}_{k}^{i}=\theta\tilde{\alpha}_{k-1}^{i}\leq\theta\Delta_{k-1}$
    or
    $\tilde{\alpha}_{k}^{i}=\theta c\Delta_{k-1}\leq\theta\Delta_{k-1}$
    since $c\in(0,1]$. Hence, by Lemma 1, we have
@@
    $\begin{split}\Phi_{k}-\Phi_{k-1}&=f(x_{k})-f(x_{k-1})+\frac{1}{2}c^{2}\gamma\left(\Delta_{k}^{2}-\Delta_{k-1}^{2}\right)\\ &\leq\frac{1}{2}c^{2}\gamma\left(\Delta_{k}^{2}-\frac{1}{\theta^{2}}\Delta_{k}^{2}\right)\\ &=-\frac{1}{2}c^{2}\gamma\left(\frac{1-\theta^{2}}{\theta^{2}}\right)\Delta_{k}^{2}.\end{split}$ (22)
@@
Using the above result, we can easily show the convergence to zero of the sequences of tentative and actual stepsizes produced by the algorithm, i.e., $\{\tilde{\alpha}_{k}^{i}\}$, $\{\alpha_{k}^{i}\}$, $i=1,\dots,n$, respectively, together with the convergence to zero of the sequence of maximum tentantive stepsizes $\{\Delta_{k}\}$.

### v3

@@ 1 Introduction
In the current paper, we analyze a line search method to solve problem (1). The algorithm under analysis is a modification of the one proposed in [29], equipped with a line search technique described in [7]. In particular, for any considered direction (i.e., a vector of the canonical basis), the line search technique first checks for a sufficient decrease in the objective function using a given stepsize. Then, if such a decrease is obtained, an extrapolation (or expansion) phase starts, where increasingly larger values of the stespize are tried until some conditions are met. This approach allows us to obtain complexity and identification results that extend those existing in the literature for direct-search [19] and model-based [22] methods. In particular, let us summarize the main contributions of this paper below.
@@
    Here, we show that the proposed algorithm correctly identifies the active constraints satisfying the strict complementarity condition in a finite number of iterations, without using any parameter-dependent estimate. Namely, this feature is just an intrinsic property of the proposed algorithm. Such a result is obtained by using some tools from the analysis of derivative-based methods  [13, 32]. More specifically, we define a measure which represents the minimum strict complementarity among the active constraints, thereby providing a neighborhood of the limit points where the active-set identification holds. Let us remark that also this identification result is obtained thanks to the extrapolation technique used in the line search procedure, which allows the stepsize to expand until we hit the border of the feasible set when we are in a neighborhood of a stationary point.
@@ 2 The algorithm
This section is concerned with the definition of a line search algorithm to solve problem (1). The proposed method, denoted as Algorithm 1, is inspired by the method proposed in  [28] and uses some ideas from the one proposed in [7] for unconstrained problems.
@@
Given a feasible point $x$, the exploration of the $i$th coordinate direction $e_{i}$ is performed by a line search procedure outlined in Algorithm 2. First, we check if the given stepsize $\nu$ is feasible along ±$\pm e_{i}$, that is, if either of $x+\nu e_{i}$ and $x-\nu e_{i}$ is feasible. If this is not the case, then we quit the line search returning a zero step length to indicate a failure. Otherwise, we try to determine if one between $e_{i}$ and $-e_{i}$ is a “good” descent direction, that is, if a sufficient decrease in the objective function can be obtained by using a feasible stepsize. If neither $e_{i}$ nor $-e_{i}$ qualifies as a suitable descent direction, then the line search procedure terminates, returning a zero step length to indicate a failure. Conversely, if a sufficient decrease of $f$ is obtained, then an extrapolation (or expansion) phase starts (i.e., lines 12–15), where we try to increase the stepsize to the maximum extent while preserving feasibility and guaranteeing the sufficient decrease condition. Specifically, the while loop keeps expanding the stepsize as long as the most recently accepted point remains strictly within the bounds (i.e., $\alpha<\alpha_{\text{max}}$) and the new tentative point is sufficiently better than the last accepted one (i.e., $f(x+\omega d)\leq f(x+\alpha d)-\gamma(\omega-\alpha)^{2}$).
@@
12:set $\alpha=\bar{\alpha}$ and $\omega=\min\{\alpha/\delta,\alpha_{\text{max}}\}$
13:while ($\alpha<\alpha_{\text{max}}$ and $f(x+\omega d)\leq f(x+\alpha d)-\gamma(\omega-\alpha)^{2}$) do
14:     set $\alpha=\omega$ and $\omega=\min\{\alpha/\delta,\alpha_{\text{max}}\}$
@@ 3.1 Assumptions and preliminary results
The above measure has been successfully used in the analysis of some direct search methods for linearly constrained problems [19, 23, 24]. It can be interpreted as the progress on a first-order model in a ball centered at $x$ with unit radius subject to feasibility constraints [11, 24], thus generalizing $\|\nabla f(x)\|$ from the unconstrained setting. Originally proposed in [10] for more general constraints and further analyzed in [11], $\chi(x)$ is continuous, non-negative and such that $\chi(x)=0$ if and only if $x$ is a KKT point. So, we can define a stationary point as follows.
@@
$\Delta_{k+1}=\tilde{\alpha}_{k+1}^{\bar{\imath}}=\theta\Delta_{k},$
which concludes the proof. ∎
@@ 3.2 Global convergence
$f(x+\omega d)\leq f(x+\alpha d)-\gamma(\omega-\alpha)^{2},$ (6)
with $\omega=\alpha/\delta=\bar{\alpha}/\delta^{j+1}$. Now, for $j$ sufficiently large, (6) contradicts Assumption 1, i.e., that $f$ is bounded from below on the feasible set. ∎
In the following, we present some results concerning the global convergence of Algorithm 1 to stationary points. More specifically, by extending some results from [7, 28], we establish a relationship between $\nabla f(x_{k})$, with specific directions $d$, and the largest tentative step length at $\Delta_{k+1}$. We will show that the bound depends, besides on the problem dimension $n$, on the Lipschitz constant $L$ and the algorithm parameters $\gamma$, $\theta$ and $\delta$.
@@
For all $k$, the result trivially holds for $d=0$. Now, consider an iteration $k$ such that $x_{k+1}\neq x_{k}$ (i.e., a successful iteration). The following cases can occur, recalling that the analysis is limited to considering directions ±$\pm e_{i}$, $i=1,\ldots,n$, belonging to $G_{k}$.
@@
    $e_{i}\in G_{k}$, $(y_{k}^{i+1})_{i}=l_{i}$ and $(y_{k}^{i})_{i}=l_{i}$. Then, $\alpha_{k}^{i}=0$ and $\tilde{\alpha}_{k+1}^{i}=\nu_{k}^{i}$. From the instructions of the line search procedure, we have that
@@
    $-e_{i}\in G_{k}$, $(y_{k}^{i+1})_{i}=l_{i}$ and $(y_{k}^{i})_{i}>l_{i}$ Then, $\tilde{\alpha}_{k+1}^{i}=\alpha_{k}^{i}>0$. From the instructions of the line search procedure, there exists $\beta_{k}^{i}\in\{0,\delta\tilde{\alpha}_{k+1}^{i}\}$ such that
@@
    Hence, by similar reasoning as in case (1a), we obtain
@@
    $-e_{i}\in G_{k}$, $(y_{k}^{i+1})_{i}=u_{i}$ and $(y_{k}^{i})_{i}=u_{i}$. Then, $\alpha_{k}^{i}=0$ and $\tilde{\alpha}_{k+1}^{i}=\nu_{k}^{i}$. Reasoning as in case (1a), with minor differences, we get
@@
    $e_{i}\in G_{k}$, $(y_{k}^{i+1})_{i}=u_{i}$ and $(y_{k}^{i})_{i}<u_{i}$. Then, $\tilde{\alpha}_{k+1}^{i}=\alpha_{k}^{i}>0$. Reasoning as in case (1b), with minor differences, we get
@@
    {±$\{\pm e_{i}\}\cap G_{k}\neq\emptyset$, $l_{i}<(y_{k}^{i+1})_{i}<u_{i}$ and $y_{k}^{i}=y_{k}^{i+1}$. Then, $\alpha_{k}^{i}=0$ and $\tilde{\alpha}_{k+1}^{i}=\nu_{k}^{i}$. From the instructions of the line search procedure, we have that
@@
    $d_{k}^{i}\in G_{k}$, $l_{i}<(y_{k}^{i+1})_{i}<u_{i}$ and $y_{k}^{i}\neq y_{k}^{i+1}$. Then, $\alpha_{k}^{i}>0$, $\tilde{\alpha}_{k+1}^{i}=\alpha_{k}^{i}$. Let us assume that $d_{k}^{i}=-e_{i}$, i.e., $y_{k}^{i+1}=y_{k}^{i}-\tilde{\alpha}_{k+1}^{i}e_{i}$ (the proof for the case $d_{k}^{i}=e_{i}$ is identical, except for minor changes). Then, from the instructions of the line search procedure, there exist $\beta_{k}^{i}$ and $\bar{\beta}_{k}^{i}$ such that $\beta_{k}^{i}\in\{0,\delta\tilde{\alpha}_{k+1}^{i}\}$, $\tilde{\alpha}_{k+1}^{i}<\bar{\beta}_{k}^{i}\leq\tilde{\alpha}_{k+1}^{i}/\delta$ and
@@
$\chi(x_{k})\leq\sqrt{n}\left(-\nabla f(x_{k})^{\top}d+\frac{M_{g}}{\theta}\Delta_{k+1}\right).$
@@
$\Phi_{k}=f(x_{k})+{\eta}\Delta_{k}^{2},$ (16)
where $\eta$ satisfies
$0<\eta<\gamma(\delta(1-\delta))^{2}.$ (17)
@@
$c_{1}=\min\left\{\gamma c^{2},\gamma(\delta(1-\delta))^{2}-\eta,\eta\left(\frac{1-\theta^{2}}{\theta^{2}}\right)\right\}>0,$ (20)
@@
  • $x_{k}\neq x_{k-1}$ (i.e., $k-1$ is a successful iteration) and $\Delta_{k}=\Delta_{k-1}$.
    Hence, there exists a coordinate ${\bar{\imath}}$ such that we have moved from $x_{k-1}$ along ±$\pm e_{\bar{\imath}}$ and we have
@@
    $\Phi_{k}-\Phi_{k-1}=f(x_{k})-f(x_{k-1})+\eta\left(\Delta_{k}^{2}-\Delta_{k-1}^{2}\right)\leq-\gamma c^{2}\Delta_{k}^{2}.$ (21)
@@
    Let $h_{k-1}^{i}$ be the number of times the stepsize related to the $i$th coordinate is expanded on iteration $k-1$, that is, if $\alpha_{k}^{i}>0$
    $\alpha_{k-1}^{i}=\min\{\alpha_{\text{max}},\delta^{-h_{k-1}^{i}}\nu_{k-1}^{i}\},$
    where $\alpha_{\text{max}}$ is the distance to the bound of the $i$th coordinate over $d_{k-1}^{i}$. Now, considering $\bar{\imath}$, we have
    $\alpha_{k-1}^{\bar{\imath}}>\Delta_{k-1}\geq\nu_{k-1}^{\bar{\imath}}.$
    In particular, the first inequality implies that $h_{k-1}^{\bar{\imath}}\geq 1$ (otherwise we would have $\alpha_{k-1}^{\bar{\imath}}\in\{0,\nu_{k-1}^{\bar{\imath}}\}$). If $h_{k-1}^{\bar{\imath}}=1$, then we have
    $\Delta_{k}=\alpha_{k-1}^{\bar{\imath}}=\min\{\alpha_{\text{max}},\delta^{-1}\nu_{k-1}^{\bar{\imath}}\}\leq\delta^{-1}\nu_{k-1}^{\bar{\imath}},$
    that is,
    $\nu_{k-1}^{\bar{\imath}}\geq\delta\Delta_{k}.$
    Thus, we get
    $f(x_{k})\leq f(x_{k-1})-\gamma(\nu_{k-1}^{\bar{\imath}})^{2}\leq f(x_{k-1})-\gamma\delta^{2}\Delta_{k}^{2}.$ (22)
    Let us now consider $h_{k-1}^{\bar{\imath}}\geq 2$. We have
    $\Delta_{k}=\alpha_{k-1}^{\bar{\imath}}=\min\{\alpha_{\text{max}},\delta^{-h_{k-1}^{\bar{\imath}}}\nu_{k-1}^{\bar{\imath}}\}\leq\delta^{-h_{k-1}^{\bar{\imath}}}\nu_{k-1}^{\bar{\imath}},$
    that is,
    $\delta^{-(h_{k-1}^{\bar{\imath}}-1)}\nu_{k-1}^{\bar{\imath}}\geq\delta\Delta_{k}.$ (23)
    Therefore, we can write
    $\begin{split}\displaystyle f(x_{k})&\leq f(x_{k-1})-\gamma\left(\delta^{-(h_{k-1}^{\bar{\imath}}-1)}\nu_{k-1}^{\bar{\imath}}-\delta^{-(h_{k-1}^{\bar{\imath}}-2)}\nu_{k-1}^{\bar{\imath}}\right)^{2}\\ &=f(x_{k-1})-\gamma\left(\delta^{-(h_{k-1}^{\bar{\imath}}-1)}\nu_{k-1}^{\bar{\imath}}(1-\delta)\right)^{2}\\ &\leq f(x_{k-1})-\gamma\left(\delta(1-\delta)\right)^{2}\Delta_{k}^{2},\end{split}$ (24)
    where the last inequality follows from (23). Finally, using (22) and (24), and considering that $\min\{(1-\delta)^{2},(\delta(1-\delta))^{2}\}=(\delta(1-\delta))^{2}$, we can write
    $f(x_{k})\leq f(x_{k-1})-\gamma(\delta(1-\delta))^{2}\Delta_{k}^{2}.$
@@
    $\begin{split}\Phi_{k}-\Phi_{k-1}&=f(x_{k})-f(x_{k-1})+\eta\left(\Delta_{k}^{2}-\Delta_{k-1}^{2}\right)\\ &\leq-\gamma(\delta(1-\delta))^{2}\Delta_{k}^{2}+\eta\Delta_{k}^{2}\\ &=-\left(\gamma(\delta(1-\delta))^{2}-\eta\right)\Delta_{k}^{2}.\end{split}$ (25)
  • $x_{k}=x_{k-1}$ (i.e., $k-1$ is an unsuccessful iteration). By Lemma 1, we have
@@
    $\begin{split}\Phi_{k}-\Phi_{k-1}&=f(x_{k})-f(x_{k-1})+\eta\left(\Delta_{k}^{2}-\Delta_{k-1}^{2}\right)\\ &=\eta\left(\Delta_{k}^{2}-\frac{1}{\theta^{2}}\Delta_{k}^{2}\right)\\ &=-\eta\left(\frac{1-\theta^{2}}{\theta^{2}}\right)\Delta_{k}^{2}.\end{split}$ (26)
@@
Using the above result, we can easily show the convergence to zero of the sequences of tentative and actual stepsizes produced by the algorithm, i.e., $\{\tilde{\alpha}_{k}^{i}\}$, $\{\alpha_{k}^{i}\}$, $i=1,\dots,n$, respectively, together with the convergence to zero of the sequence of maximum tentative stepsizes $\{\Delta_{k}\}$.